OpenStudy (anonymous):

A motorboat can maintain a constant speed of 16 miles per hour relative to the water. The boat makes a trip upstream to a certain point in 20 minutes; the return trip takes 15 minutes. What is the speed of the current.

OpenStudy (anonymous):

Let x = speed of the current then (16-x) = speed upstream and (16+x) = speed downstream : Convert 20 min to hrs 20%2F60 = 1%2F3 hrs Convert 15 min to hrs 15%2F60 = 1%2F4 hrs : Distance of the two trips is the same; write a dist equation: dist = time * speed ; 1%2F4(16+x) = 1%2F3(16-x) Multiply both sides 12 to get rid of the denominators: 3(16 + x) = 4(16-x) : 48 + 3x = 64 - 4x : 3x + 4x = 64 - 48 : 7x = 16 x = 16%2F7 x ~ 2.3 mph is the current : Check solution with a calc, confirm that the distances are equal .33*13.7 = 4.5 .25*18.3 = 4.5 :